A scientist has two solutions,which she has labeled Solution A and B.
Each contains salt. She knows that Solution A is 65% SALT and Solution B is 90%
SALT. She wants to obtain 120 ounces of a mixture that is 85% salt.
How many ounces of EACH solution should she use?
Solution A: _ ounces
Solution B:_ ounces
Thanks.

Found 2 solutions by galactus, modi.kk91:
Answer by galactus(183)

Let A=amount of solution A
.65A+.90(120-A)=.85(120)
Solve for A and note that amount of B is 120-A

Answer by modi.kk91(20)

by allegation mixture rule

65% 90%

85%



90-85=5 and 85-65=20


ratio=5:20=1:4=====1x+4x=5x
ans==1x=24
4x=96
5x=120
x=24